40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 19 09:36:32 2013, Koji, Update, Green Locking, Xend green aligned ALS_OUTOFLOOP_130816.pdf
    Reply  Mon Aug 19 14:22:36 2013, rana, Update, Green Locking, Xend green aligned 
       Reply  Mon Aug 19 17:40:32 2013, Steve, Update, Green Locking, Xend green layout corrections beforeC.jpgnowC.jpgstillMultiple.jpg
          Reply  Mon Aug 19 19:08:35 2013, Koji, Update, Green Locking, Xend green layout corrections 
          Reply  Tue Aug 20 10:59:15 2013, Steve, Update, Green Locking, Xend green layout corrections faraway.jpgcloser.jpg
             Reply  Wed Aug 28 11:15:02 2013, Steve, Update, Green Locking, Xend green layout corrections 
    Reply  Mon Aug 19 16:18:56 2013, manasa, Update, Green Locking, Xend green aligned 
Message ID: 9083     Entry time: Wed Aug 28 11:15:02 2013     In reply to: 9039
Author: Steve 
Type: Update 
Category: Green Locking 
Subject: Xend green layout corrections 

Quote:

Quote:

 Shutter moved, no more clipping.

Pick-off mirror 2" replaced by 1" one. Laseroptik HR 532nm, incident angle 30-45 degrees, AR 532 nm

Green REFL PD moved to 4" close to pick-off mirror. Pd being close to pick-off does not separate multiple reflections on it. I'll replace Laseroptic mirror with Al one. It is not easy to find.

 Hole cut into side wall for doubler oven cable to exit.

 

 

 Beam trap for Pd refl is in place. Cabeling is ti·died up.

 Laseroptic 1" mirror is replaced by Al 1" mirror. Problem remains the same. This diffraction patter has to be coming from the Faraday.

  Atm1, good separation when Pd is far 

  Atm2, bad separation when Pd is close 

 The extra high post 3.375"  for PZT is ready. We also have 2 more 2" green Laseroptik mirrors. I'm ready to swap them in.

The  75 mm  focal length  lens  was placed in front of the green REFL PD yesterday.

ELOG V3.1.3-